OpenStudy (anonymous):

Consider the function f(x) = 3x-8, What is the inverse of f(x) and name it g(x). I got that the answer was f^1(x) 1/3x+8/3. My teacher won't take though unless I show my work. Someone please help? I'm desperate?

OpenStudy (cwrw238):

this is the way to find the inverse let y = 3x -8 now find x in terms of y: 3x = y - 8 x = (y-8) / 3 now repalce y by x ;- inverse of f(x) = g(x) = (x - 8)/ 3

OpenStudy (anonymous):

Was I correct though?

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

oopp sorry i made a mistake!!! it should be ( x + 8) / 3

OpenStudy (cwrw238):

or x/3 +8/3 - i've been making typos all day !!

OpenStudy (anonymous):

so which is it...?

OpenStudy (cwrw238):

can u understand how i got the answer or do you need more detail

OpenStudy (anonymous):

I need more detail. my answer wasn't too far from the one you got, But I got 1/3 +8/3...

OpenStudy (cwrw238):

ok let y = 3x - 8 add 8 to both sides y + 8 = 3x now divide both sides by 3 (y + 8) / 3 = x flipping the sides x = (y + 8/ 3 f-1 (x) = (x + 8)/ 3 - this is the correct answer another way of writing it is x/3 + 8/3

OpenStudy (cwrw238):

each of these answers is correct - take your pick

OpenStudy (anonymous):

How would I use function composition to show that f(x) and gx) are inverses of eachother?

OpenStudy (cwrw238):

ok if f(x) and g(x) are inverses of eavh other then f(g(x)) = x lets check :- f(g(x) = 3 (x+ 8) ----- - 8 3 = 3x + 24 - 24 -------------- 3 = 3x /3 = x g(f(x)) will also be = x